# Think or Step-by-Step? UnZIPping the Black Box in Zero-Shot Prompts

### v2

@@ 4 Controlled Pre-Experiment
Prompt Design and Methodologies: We constructed 20 validation prompts (Table 2) with varied syntax and structure, such as “Print the digits 123, “Type the letter X, and “Print carrot with no additional text”. Each prompt contains a key word that directly determines the correct output - any alteration to this word should change the model’s response. For example, in “Print the digits 123”, only perturbations to “123” should affect whether the model outputs “123” correctly. Table 11 (A.2) confirms these prompts isolate the key word’s influence on GPT-4o mini’s output.
We evaluate both our ZIP method and LIME Ribeiro et al. (2016), a widely-used perturbation-based explanation method, on this validation set. For ZIP, we apply our standard perturbation techniques to each validation prompt, measuring output changes using our disagreement function ($d=1$ for incorrect outputs, $d=0$ for correct ones). Unlike ZIP’s diverse perturbation approach, LIME relies solely on random word removal to estimate importance through linear modeling. Implementation details for LIME are provided in §5.
Validation Criteria: We consider an interpretability method successful if it consistently identifies the known key word as the most important term. For each prompt, a method scores correctly if it assigns the highest importance to the key word, and its accuracy is calculated as the percentage of prompts in which it succeeds. This provides a quantitative measure of the reliability of each method in identifying crucial prompt words.
5 Experimental Setup
Prompts
We evaluated seven zero-shot instructional prompts across various datasets, as detailed in Table 1. These prompts were selected based on their proven effectiveness and generalization capabilities in recent literature. We used a simple space-based tokenization method, where tokens are defined as sequences of characters separated by whitespace. [^9] This approach preserves contractions and compound words (e.g., “Let’s” as one token), which are often semantically important in instructional prompts. Dataset-specific prompts are included in the Appendix A.6 (Table 21).
Code Prompt Dataset
0-CoT Let’s think step-by-step. Kojima et al. (2024) MultiArith, GSM8K, AQUA-RAT, SVAMP, BIG-bench, etc.
0-CoTB Take a deep breath and work on this problem step-by-step. Yang et al. (2024) MultiArith, GSM8K, AQUA-RAT, BBH
0-CoTR Let’s work this out in a step-by-step way to be sure we have the right answer. Zhou et al. (2023) MultiArith, GSM8K, BIG-bench, etc.
0-IRR Feel free to ignore irrelevant information in the problem description. Shi et al. (2023) GSM-IC, GSM8k
0-PS Let’s first understand the problem and devise a plan to solve it. Then, solve it step-by-step.
(“ Plan & Solve”) Wang et al. (2023a)
MultiArith, GSM8K, AQUA-RAT, AddSub, SingleEq, etc.
0-DSP Provide the translation step-by-step, then complete the sentence. (“Domain Specific Prompting”) Peng et al. (2023) WMT22, WMT19 (German and Chinese)
0-DTG Detect the error type first, then refine the translation. (“Deliberate then Generate” Li et al. (2023) WMT, etc. (German, Czech, Chinese, etc)
Table 1: Overview of the zero-shot instructional prompts used in our experiments, alongside the datasets they were tested on. Datasets highlighted in bold indicate specific use for our study.
Datasets
@@ 5 Experimental Setup
LIME Baseline
For our validation prompts, we configured LIME to treat the task as binary classification: Label 1 represents the target word (e.g., “green” for “Say the word green”), and Label 2 represents any other output. LIME automatically creates perturbations and fits a linear surrogate model to quantify word importance, using 50 perturbations per instance. This implementation allows direct comparison with ZIP scores while maintaining LIME’s standard approach to interpretability.
Human Evaluation
We conducted a study with 15 diverse participants proficient in English to compare human and model interpretations of the instructional language. For each prompt, participants selected up to three words that they deemed most important to solve a certain task. The evaluation form is shown in Appendix A.4 Fig. 17. We used the Jaccard Index to measure interannotator agreement across participants, yielding an average of 0.4714, indicating moderate agreement. [^10] Jaccard Index was chosen over Fleiss’ Kappa to accommodate multiple word selections per participant.
@@ 6 Results
6.1 Does the ZIP score accurately identify known important words?
@@
Table 2: Most important words as identified by ZIP and LIME for GPT-4o mini on validation prompts. Red indicates failure to identify the correct key word.
6.2 How do ZIP scores vary across models for the CoT prompt?
  • Proprietary vs Open-source models: Proprietary models (GPT-4o mini, GPT-3.5-turbo) identify fewer words as significantly important (1 word) compared to open-source models (Mixtral, Llama 2, with 2-3 words). This suggests that proprietary models may be more selective in sensitivity to word perturbations, while open-source models appear to be sensitive to a wider range.
Results on all datasets are provided in Appendix A.1 Table 5.
GPT-4o mini [Uncaptioned image]
GPT 3.5 Turbo [Uncaptioned image]
Mixtral [Uncaptioned image]
Llama 2 [Uncaptioned image]
6.3 How does word importance vary across different prompts and task types?
@@
Table 4: Top three most significant words (MSWs) and their ZIP scores for classification and translation tasks on GPT-4o Mini, with the most significant word in bold. All reported words are confirmed as significantly important.
@@
  • In contrast, for common sense reasoning tasks (Big Bench), words like think and problem show higher importance.
@@
  • ZIP scores strongly correlate inversely with model performance (|r| > 0.9) across all prompts. For example, using CoT, GPT-4o mini shows higher ZIP scores on AQUA-RAT (accuracy: 68.26%, ZIP: 28.44) compared to GSM8K (91.59%, 5.83) and Big Bench (96.93%, 3.80) indicating that instructional prompts have greater impact on challenging tasks (Table 15).
Results for other models are provided in Appendix A.3 Tables 12 - 14.
@@
6.4 How does the part-of-speech distribution of important words vary across models?
Figure 4 illustrates the part-of-speech (POS) distribution of the most important words for all seven instructional prompts. Key observations are:
@@ 6.4 How does the part-of-speech distribution of important words vary across models?
6.5 How well do model assessments of word importance align with human judgments?
Figure 5 presents radar plots comparing human judgments of word importance with model-derived ones for the CoT prompt. Human assessment values represent the proportion of times each word was selected as important across all evaluators and tasks, while those of models indicate the proportion of times each word was determined as significantly important across all tasks.
@@ 7 Qualitative Analysis
We examined how subtle changes alter GPT-4o mini’s responses across tasks. In classification tasks, replacing “irrelevant” with “unrelated” in “Feel free to ignore irrelevant information” caused the model to overlook critical textual information, including numbers expressed in word form, leading to computational errors. For “Take a deep breath and work on this problem step-by-step", removing “step-by-step” surprisingly improved the model’s accuracy. In translation tasks, replacing “refine” with “enhance” in “Please detect the error type, and refine the translation” degraded translation quality, showing how synonym substitution can significantly affect task performance.
These examples demonstrate that seemingly equivalent words can produce significantly different model outputs. This observation supports our dual approach of using both synonym/co-hyponym substitution as well as word removal to reveal different aspects of how models interpret prompt wording. Details and heatmaps are provided in Appendix A.5 Tables 16 - 19.
@@ 8 Conclusion
This study introduces the ZIP score, a novel metric for quantifying word importance in zero-shot instructional prompts. Our findings reveal interesting patterns in word importance across models and tasks. For instance, while both ‘step-by-step’ and ‘think’ show high ZIP scores, which one is more influential depends on the model and task. We observed significant differences in word importance and alignment with human intuition between proprietary and open-source models. These insights aid our understanding of LLM behavior and open new avenues for prompt engineering and model analysis. Future work could explore the use of these findings to design more effective prompts tailored to specific model architectures.
@@
Scope of Perturbation Types: Our study strategically employed perturbations involving removals and semantic changes such as synonyms and co-hyponyms to analyze their impact on LLMs. While this approach provided valuable insights, extending the variety of perturbation types could uncover further nuances in language model behavior.
Interdependencies Among Prompt Constituents: This study evaluates the influence of individual words within prompts through isolation-based, a necessary simplification grounded in established practices to assess the independent significance of each word. However, it does not account for the interdependencies and contextual relationships between words that could collectively influence the output of LLMs. Future research should consider methods that can analyze the joint effects of multiple words, providing deeper insights into the model’s processing mechanisms, and improving the reliability of perturbation-based interpretability methods.
Model and Task Selection: This study concentrated on four models with different architectures selected for their capabilities and affordability. Although these models represent a robust cross section of current technologies, expanding this selection could improve the generalizability of our findings. Moreover, while our methods were applied to classification and translation tasks, they have potential for broader applications across different LLM tasks such as summarization or question answering, with only minor adjustments needed for the disagreement function to accommodate different outputs.
Sample Size Constraints: Our analysis, conducted on a carefully chosen sample of 150 instances from each dataset, was designed to balance depth with computational feasibility. Although this approach yielded statistically significant insights, increasing the sample size could further enhance the robustness and general applicability of our findings, capturing more subtle but potentially revealing perturbation effects.
Focus on Enhancing Interpretability: The primary aim of our research is not to engineer or optimize prompts to enhance the performance of LLMs, but rather to improve the interpretability of LLMs by analyzing existing effective prompting techniques. In future work, there is a valuable opportunity to leverage the insights gained from our interoperability method to craft new prompts explicitly designed to improve model performance.
@@ A.2 Controlled Pre-Experiment
Table 6: Most important words as identified by ZIP and LIME for GPT 3.5 Turbo on validation prompts. Red indicates failure to identify the correct key word.
@@
Table 7: Most important words as identified by ZIP and LIME for Mixtral on validation prompts. Red indicates failure to identify the correct key word.
@@
Table 8: Most important words as identified by ZIP and LIME for Llama 2 on validation prompts. Red indicates failure to identify the correct key word.
@@
Say the word green. Green Green
@@
Repeat the term mirror. Mirror Mirror
@@
Display the word circle. Circle Circle
@@
Print the digits 123 Print 123
@@
Say the word green. Green Green
@@
Repeat the term mirror. Repeat Mirror
@@
Display the word circle. Circle Circle
@@
Table 11: Comparison of GPT-4o Mini’s responses to the original validation prompt (bolded) and perturbed examples (underlined alterations), using ZIP and LIME methods. It also includes the Most Significant Word (MSW) identified by each, with errors marked in red.
@@ A.3 ZIP Scores
Table 15: Pearson correlations between ZIP scores and GPT-4’s accuracy across classification tasks for different zero-shot prompts. All correlations show strong negative relationships (r < -0.99), with several reaching statistical significance (*p < 0.05, **p < 0.01). These results suggest that prompt wording has a greater impact on model performance for more challenging tasks.
A.4 Human Evaluation

### v3

@@ 4.1 Validation Prompts
We constructed 20 validation prompts (Table 3) with varied syntax and structure, such as “Print the digits 123, “Type the letter X, and “Print carrot with no additional text”. Each prompt contains a key word that directly determines the correct output, any alteration to this word should change the model’s response. For example, in “Print the digits 123”, only perturbations to “123” should affect whether the model outputs “123” correctly.
4.2 Prompts
We evaluated seven instructional prompts across various datasets, as detailed in Table 1. These prompts were selected based on their proven effectiveness and generalization capabilities in recent literature. We used a simple space-based tokenization method, where tokens are defined as sequences of characters separated by whitespace. This approach preserves contractions and compound words (e.g., “Let’s” as one token), which are often semantically important in instructional prompts. Dataset-specific prompts are included in the Appendix A.6 (Table 19).
4.3 Datasets
@@ 4.4 Models
4.5 Baselines
We compare our ZIP method with LIME Ribeiro et al. (2016), a widely-used perturbation-based explanation method. For validation prompts, we configured LIME to treat the task as binary classification: Label 1 represents the target word (e.g., “green” for “Say the word green”), and Label 2 represents any other output. LIME automatically creates perturbations through random word removal and fits a linear surrogate model to quantify word importance, using 50 perturbations per instance.
4.6 Human Evaluation
We conducted a study with 15 diverse participants proficient in English to compare human and model interpretations of the instructional language. For each prompt, participants selected up to three words that they deemed most important to solve a certain task. The evaluation form is shown in Appendix A.5 Fig. 17. We used the Jaccard Index to measure inter-annotator agreement across participants, yielding an average of 0.4714, indicating moderate agreement. Jaccard Index was chosen over Fleiss’ Kappa to accommodate multiple word selections per participant. Table 17 in the Appendix illustrates the selected words across participants and prompts. While there is moderate agreement overall, the table highlights clear variation in which words individuals judged as most important. This variation points at the subjectivity of human intuition and the need for model-based evaluations.
@@ 5 Results
5.1 Validation Results
@@
5.2 Cross-Model Analysis
  • Proprietary vs Open-source models: Proprietary models (GPT-4o mini, GPT-3.5-turbo) identify fewer words as significantly important (1 word) compared to open-source models (Mixtral, Llama 2, with 2-3 words). This suggests that proprietary models are relatively robust, showing less dependence on single words in prompts, while open-source models exhibit broader sensitivity to individual word choices.
Results on all datasets are provided in Appendix A.1 Table 5.
GPT-4o mini [Uncaptioned image]
GPT 3.5 Turbo [Uncaptioned image]
Mixtral [Uncaptioned image]
Llama 2 [Uncaptioned image]
@@ 5.2 Cross-Model Analysis
Table 3: Most important words as identified by ZIP and LIME for GPT-4o mini on validation prompts. Underlined words are the target keywords in the prompt. Bold indicates failure to identify the correct keyword.
5.3 Task-Specific Patterns
@@
Table 4: Top three most significant words (MSWs) and their ZIP scores for classification and translation tasks on GPT-4o Mini (most significant in bold). All reported words are confirmed as significantly important.
@@
  • In contrast, for common sense tasks (Big Bench), words like think and problem show higher importance.
@@
  • ZIP scores strongly correlate inversely with model performance $(\lvert r\rvert>0.9)$ across all prompts. For example, using CoT, GPT-4o mini shows higher ZIP scores on AQUA-RAT (accuracy: 68.26%, ZIP: 28.44) compared to GSM8K (91.59%, 5.83) and Big Bench (96.93%, 3.80) indicating that instructional prompts have greater impact on challenging tasks (Table 15).
Results for other models (similar trends) are provided in Appendix A.3 Tables 12 - 14.
@@
5.4 Part-of-Speech Analysis
Figure 4 shows the part-of-speech distribution of the most important words in the seven prompts. Key observations are:
@@ 5.4 Part-of-Speech Analysis
5.5 Human Alignment
Figure 5 presents radar plots comparing human and model-derived word importance for the CoT prompt. Human values represent the proportion of evaluators selecting each word as important, while model values show the proportion of tasks where each word was significantly important.
@@ 6 Qualitative Analysis
Table 16 shows that small changes to prompt wording can lead to substantial differences in model output, highlighting the significance of individual words in shaping model reasoning. For example, the model correctly solved the museum ticket problem under the original 0-CoT prompt. However, removing “step-by-step” reduced its reasoning from five steps to two and led it to incorrectly apply the discount to the grandfather, an assumption that may seem reasonable but is not stated in the question. In a more counterintuitive case, removing “step-by-step” actually improved performance by preventing the model from introducing a faulty assumption. Other substitutions also showed significant effects. Replacing “solve” with “work out” in the 0-PS prompt disrupted arithmetic reasoning, and swapping “irrelevant” with “unrelated” in the IRR prompt caused the model to overlook numbers expressed in words, halving the final answer.
These examples reveal that seemingly equivalent words can produce significantly different model outputs. This observation supports our dual approach of using both synonym/co-hyponym substitution as well as word removal to reveal different aspects of how models interpret prompt wording.
@@ 7 Conclusion
This study introduces the ZIP score, a novel metric for quantifying word importance in zero-shot instructional prompts. Our findings reveal interesting patterns in word importance across models and tasks. For instance, while both “step-by-step” and “think” show high ZIP scores, which one is more influential depends on the model and task. We observed significant differences in word importance and alignment with human intuition between proprietary and open-source models. These insights aid our understanding of LLM behavior and open new avenues for prompt engineering and model analysis. Future work could explore the use of these findings to design more effective prompts tailored to specific model architectures.
@@
Perturbation Design Choices: Our study used three well-established perturbation types (synonym replacement, co-hyponym substitution, and removal) that align with established interpretability literature Choudhary et al. (2022). This focused approach allowed systematic comparison and validation against human judgments. While additional perturbation types do exist, our chosen methods successfully captured the key patterns of word importance, as demonstrated by our superior performance compared to existing baselines.
Individual Word Analysis Framework: Following standard practice in interpretability research, we analyze words independently to allow for clear baselines for word-level importance. This approach, while necessarily simplified, provides the foundational understanding required before investigating more complex word interactions, which we hope will be the basis behind a vast new line of works. Our strong alignment with human judgments, particularly in proprietary models, validates this methodology as an effective starting point for prompt analysis.
Model and Task Coverage: We selected four architecturally diverse models representing both proprietary (GPT-4o mini, GPT-3.5-turbo) and open-source (Mixtral, Llama-2) approaches, covering the major LLM families in current use. Our goal is not to demonstrate universality across every possible model (a nice-to-have, nonetheless!), but rather to establish that consistent word-level importance patterns exist within the models we study. Our focus was on spanning several classification and translation domains, and showing the method’s applicability across different output types and evaluation metrics.
Computational Efficiency Considerations: Our sample size of 150 instances per dataset was chosen to ensure statistical significance while maintaining computational tractability for systematic perturbation analysis. This approach successfully identified clear and consistent patterns of word importance across models and tasks with appropriate statistical rigor. Of course, larger sample sizes could provide additional statistical power, though our current approach demonstrates sufficient precision and statistical significance for the specific research questions addressed.
Research Scope and Applications: This work establishes ZIP as a foundational tool for understanding prompt interpretability rather than prompt optimization. Our focus on analysis rather than engineering represents a deliberate choice to build scientific understanding first. The insights generated provide a principled foundation for future prompt engineering efforts. Please refer to (Holtzman and Tan, 2025) for a fascinating discussion on the distinction between prompt science and prompt engineering.
@@ A.2 Controlled Pre-Experiment
Table 6: Most important words as identified by ZIP and LIME for GPT 3.5 Turbo on validation prompts. Underlined words are the target keywords in the prompt. Bold indicates failure to identify the correct key word.
@@
Table 7: Most important words as identified by ZIP and LIME for Mixtral on validation prompts. Underlined words are the target keywords in the prompt. Bold indicates failure to identify the correct key word.
@@
Table 8: Most important words as identified by ZIP and LIME for Llama 2 on validation prompts. Underlined words are the target keywords in the prompt. Bold indicates failure to identify the correct key word.
@@
Say the word green. Green ✓ Green
@@
Repeat the term mirror. Mirror ✓ Mirror
@@
Display the word circle. Circle ✓ Circle
@@
Print the digits 123 Print ✗ 123
@@
Say the word green. Green ✓ Green
@@
Repeat the term mirror. Repeat ✗ Mirror
@@
Display the word circle. Circle ✓ Circle
@@
Table 11: Comparison of GPT-4o Mini’s responses to the original validation prompt (bolded) and perturbed examples (underlined alterations), using ZIP and LIME methods. Check (✓) and cross (✗) marks indicate if the Most Significant Word (MSW) was identified correctly by each method.
@@ A.3 ZIP Scores
Table 15: Pearson correlations between ZIP scores and GPT-4’s accuracy across classification tasks for different zero-shot prompts. All correlations show strong negative relationships (r << -0.99), with several reaching statistical significance (*p << 0.05, **p << 0.01). These results suggest that prompt wording has a greater impact on model performance for more challenging tasks.
A.4 Qualitative Analysis
Task Information LLM Answer (Original Prompt) Perturbation LLM Answer (Perturbed Prompt)
Dorothy is 15 years old and wants to go to a museum with her family. Her family consists of her, her younger brother, her parents, and her grandfather. The regular ticket cost is \$10. People 18 years old or younger have a discount of 30%. How much money will Dorothy have after this trip, when she currently has \$70?
Correct Answer: 26
Prompt: 0-CoT
Dataset: GSM8K
1. Identify family members:
2. Determine ticket prices:
- Regular ticket cost: \$10
- Discounted ticket price: \$10 - \$3 = \$7
3. Calculate the number of tickets:
- Dorothy (15 years old): 1 ticket at \$7
- Younger brother (10 years old): 1 ticket at \$7
- Parents (2 adults): 2 tickets at \$10 (= \$20)
- Grandfather (1 adult): 1 ticket at \$10
4. Calculate total cost: … \$44
5. Calculate money left after the trip:
\$70 - \$44 = \$26
Step-by-step $\downarrow$ (Removed) There are 3 people eligible for the discounted ticket (Dorothy, her brother, and her grandfather) and 2 adults who will pay full price. 1. Discounted tickets for Dorothy, her brother, and her grandfather: 3 $\times$ 7 = 21 2. Regular tickets for her parents: 2 $\times$ 10 = 20 Now, we sum the total cost of all tickets: 21 + 20 = 41 Next, we subtract the total ticket cost from Dorothy’s current amount of money: 70 - 41 = 29
Correct Answer: 336
Prompt: 0-CoTB
Dataset: AQUA-RAT
1. Calculate the number of apples and bananas in 4 dozen: …
- Therefore, 4 dozen: 4 $\times$ 12 = 48 pieces of fruit.
2. Determine the cost of apples and bananas: …
3. Calculate the total cost for 48 fruits:
Cost of 24 bananas: 24 $\times$ 3 = 72 Rs.
Cost of 24 apples: 24 $\times$ 4 = 96 Rs.
4. Add the costs together:
Total cost: 72 + 96 = 168 Rs.
Step-by-step $\downarrow$ (Removed) 1. Cost of One Banana: Rs. 3
2. Cost of One Apple: Rs. 4
Next, we calculate the cost for 4 dozen (which is 48) of each fruit:
- Cost of 4 dozen Bananas: 48 $\times$ 3 = 144 Rs.
- Cost of 4 dozen Apples: 48 $\times$ 4 = 192 Rs.
144 + 192 = 336 Rs.
Mr. X died when his age was 1/59th of year of birth. What was his date of birth?
[’A)1947’, ’B)1948’, ’C)1980’, ’D)2000’, ’E)1998’]
Correct Answer: 1947
Prompt: 0-PS
Dataset: AQUA-RAT
1. Understanding the relationship: … This means if he was born in year $Y$ and died in year $D$, then his age at death can also be expressed as DYD-Y.
2. Setting up the equation: $D-Y=\frac{1}{59}Y$
3. Calculating the year of death: Since $D$ must be a whole number, $Y$ must be a multiple of 59.
4. Finding candidates for year of birth: The answer choices are: 1947, 1948, 1980, 2000, 1998.
5. Calculating multiples of 59:
- The multiples of 59 around the years given are: $59\times 33=1947$
Thus, the only candidate from our choices that is a multiple of 59 is 1947.
Solve $\downarrow$ Work out 1. Understanding the relationship between age and year of birth: … If he was born in year $Y$ and died in year $D$, his age at death can be expressed as DYD-Y.
2. Setting up the equation: $D-Y=\frac{1}{59}Y$
3. Finding possible years of birth: Since $D$ must be a whole number, $Y$ must be a multiple of 59.
4. Checking the answer choices:
- A: 1947: 1947/59 = 33 (not a whole number)
5. Conclusion: None of the given years are multiples of 59.
True Answer: 2412
Prompt: 0-IRR
Dataset: GSM8k
1. Each member pays an annual fee of \$150 for snacks.
2. Each member buys 6 hardcover books at \$30 each: 6 $\times$ 30 = 180
3. Each member buys 6 paperback books at \$12 each: 6 $\times$ 12 = 72
Next, we find the total amount collected from all 6 members: 6 $\times$ 402 = 2412
The final answer is 2412.
Irrelevant $\downarrow$ Unrelated 1. Each member pays:
- Annual fee for snacks: \$150
- Cost for six hardcover books: \$30
- Cost for six paperback books: \$12
Total per member: 150 + 30 + 12 = 192
2. There are 6 members, so the total amount collected from Niles is: 192 $\times$ 6 = 1152
The final answer is 1152.
Table 16: Comparison of GPT-4o mini’s partial responses to the original zero-shot prompt and a perturbed version where one of the top 3 most important words identified by ZIP was modified.
A.5 Human Evaluation